# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 6.2. Parent/child license differences
apache-2.0 $\rightarrow$ apache-2.0 |||| mit 28 0.04%
@@
apache-2.0 |||| bsd-3-clause $\rightarrow$ bsd-3-clause 8 0.01%
@@
apache-2.0 cc-by-4.0 782 2.15%
Our dataset contains 274,104 distinct parent/child relationships. We observed 66,460 instances (24%) where the licensing of a child model (i.e., derivative model) was different from that of its parent (i.e., base model). Table 15 provides an overview of the types of differences observed. In nearly a third of cases, child models specified no licensing information despite information being available for their parent(s), leading to a potential compliance violation. We also observe that once license information has been dropped, it is unlikely to be restored by later links in the chain, with license restoration behavior observed in only 15.9% of instances. A complete shift in licensing between parent and child was seen in 54.8% of cases (i.e., no licenses were shared between the two). Of these shifts, 18.5% involved the “unknown” or “other” license category. Table 16 shows the top ten such shifts, and Figure 8 provides more information on how licensing decisions shifted across class boundaries between parents and their children. While ML-specific licenses seem a good fit for models, model owners may be opting for licenses they are more familiar, and comfortable with (OSS for software developers and CC for data scientists). We observe that most shifts occur across the OSS and CC boundary (68.3%), perhaps suggesting some tension between the licensing preferences of those with data science and software development backgrounds. Additionally, depending on the specifics of the situation and the licenses involved, these shifts can also potentially introduce license incompatibilities, such as dropping non-commercial requirements or failing to apply the requirements of a copyleft license.
Figure 8. Parent/child license differences across class boundaries (Total differences: 29,708)
@@ 6.3. Dataset license and model license combinations
The licenses of datasets that are used during training can have an impact on the licensing decisions for the final ML model. Here we consider the pairings of dataset licenses and the resultant model licenses. There are 623 distinct model/dataset license combinations within our dataset across 43,455 such pairings. (The top ten most frequent can be found in Table 17.) However, since we (and Hugging Face) aggregated all “other” licenses and treated them as one quantity, these numbers likely overestimate the consistency in the space. In 41 of the 623 combinations, the license of the model exactly matches the license of at least one dataset it was trained on. The most common combination is a model licensed under the apache-2.0 license and a dataset under a custom or “other” license. In total, 11,731 (27%) pairs involve a dataset under a custom or “other” license. This can make using these datasets and models problematic since retrieving and evaluating the licensing terms can be difficult and time-consuming.
6.4. Other licensing findings
@@
A few models (188) and datasets (104) were released under more than one license. Multi-licensing is an existing phenomenon in the world of open-source software but is made more complex in the ML model context by the presence of novel license combinations that are not yet well understood. Examples include the 48 models we observed multi-licensed under apache-2.0 and cc-by-nc-4.0 as well as models under both OSS and ML licenses, such as MIT and OpenRAIL. We observe examples of such multilicensing in the most declared datasets in Table 9. We also observed datasets that were released under as many as six distinct licenses. The Hugging Face metadata provides no way of determining the relationship among these licenses, which could be either an AND or an OR. This distinction is critical to determining compliance. For example, if a user who uploaded a llama2 derivative model chose to make it available with an OR relationship between the apache-2.0 and llama2 licenses, this would contradict the exclusive llama2 licensing of its base model, resulting in a violation that could propagate to models further down the chain. While we observed only a few instances of multi-licensing, such activity motivates the need to understand the interplay between the various license classes and for Hugging Face to supply a standardized way to specify the intended relationship between licenses.
@@
Some licenses, like Llama3, impose a generational limitation (Laurent, 2004) that specifies naming requirements for derivative works. According to the license, models that are built on a model under this license must have names prefixed with “llama3” (Meta, 2024). In our dataset, there are 3,335 models licensed under llama3. However, of these, only 473 (14.2%) models correctly have names beginning with “llama3” as required by the license. Only 768 (23%) models even contain the term “llama3” in their name. Just 2,421 (72.6%) models contain the term “llama” at all, and 384 (11.5%) only contain the string “l3.” The problem is likely even worse in practice, as these statistics only include models that documented that they were under the llama3 license.

### v2

@@ 7.3. Parent/child license differences
apache-2.0 $\rightarrow$ apache-2.0 $|$ mit 28 0.04%
@@
apache-2.0 $|$ bsd-3-clause $\rightarrow$ bsd-3-clause 8 0.01%
@@
apache-2.0 $c$-by-4.0 782 2.15%
Our dataset contains 274,104 distinct parent/child relationships. We observed 66,460 instances (24%) where the licensing of a child model (i.e., derivative model) was different from that of its parent (i.e., base model). Table 19 provides an overview of the types of differences observed. In nearly a third of cases, child models specified no licensing information despite information being available for their parent(s), leading to a potential compliance issue. We also observe that once license information has been dropped, it is unlikely to be restored by later links in the chain, with license restoration behavior observed in only 15.9% of instances. A complete shift in licensing between parent and child was seen in 54.8% of cases (i.e., no licenses were shared between the two). Of these shifts, 18.5% involved the “Unknown” or “Other” license category. Table 20 shows the top ten such shifts, and Figure 3 provides more information on how licensing decisions shifted across categories between parents and their children (e.g., a difference from cc-by-4.0 to MIT), as well as how licensing decisions shifted within categories (e.g., a difference from MIT to Apache-2.0)While ML-specific licenses seem a good fit for models, model owners may be opting for licenses they are more familiar and comfortable with (OSS for software developers and CC for data scientists). We observe that most shifts occur between the OSS and CC categories (68.3%), perhaps suggesting some tension between the licensing preferences of those with data science and software development backgrounds. Additionally, depending on the specifics of the situation and the licenses involved, these shifts can also potentially introduce license incompatibilities, such as dropping non-commercial requirements or failing to apply the requirements of a copyleft license.
Figure 3. Parent/child license differences across categories (Total differences: 29,708)
@@ 7.4. Dataset license and model license combinations
The licenses of datasets that are used during training can have an impact on the licensing decisions for the final ML model. Here we consider the pairings of dataset licenses and the resultant model licenses. There are 623 distinct model/dataset license combinations within our dataset across 43,455 such pairings; however, since we (and Hugging Face) aggregated all “Other” licenses under one label, these numbers likely overestimate the consistency in the space. (The top ten most frequent can be found in Table 21.) The license of the model exactly matches the license of at least one dataset it was trained on in only 41 of the 623 combinations. The most common combination is a model licensed under the apache-2.0 license and a dataset under a custom or “Other” license, with 11,731 (27%) pairs involving a dataset under a custom or “Other” license. This can make using these datasets and models problematic since retrieving and evaluating the licensing terms can be difficult and time-consuming.
7.5. Additional licensing findings
@@
A few models (188) and datasets (104) were released under more than one license. Multi-licensing is an existing phenomenon in the world of open-source software but is made more complex in the ML model context by the presence of novel license combinations that are not yet well understood. Examples include the 48 models we observed multi-licensed under apache-2.0 and cc-by-nc-4.0 as well as models under both OSS and ML licenses, such as MIT and OpenRAIL. We observe examples of such multilicensing in the most declared datasets in Table 9. We also observed datasets that were released under as many as six distinct licenses. The Hugging Face metadata provides no way of determining the relationship among these licenses, which may be an AND or an OR, or even indicate, as observed in Sections 8.4.1 and 9.3, that portions of the dataset are under different licenses. Regardless, this distinction is critical to determining compliance. For example, if a user who uploaded a llama2 derivative model chose to make it available with an OR relationship between the apache-2.0 and llama2 licenses, this would contradict the exclusive llama2 licensing of its base model, resulting in a violation that could propagate to models further down the chain. While we observed only a few instances of multi-licensing, such activity motivates the need to understand the interplay between the various license classes and for Hugging Face to supply a standardized way to specify the intended relationship between licenses.
As a case study, we consider the most common example of multi-licensing found across models: a combination of the apache-2.0 and cc-by-nc-4.0 licenses. Depending on the relationship between these two licenses and the reasoning for including them, the combination could represent potential compliance challenges. To explore this, we conduct a manual analysis of mradermacher/ NSFW_DPO_Noromaid-7b-Mistral-7B-Instruct-v0.1-GGUF, the model under this licensing scheme with the most downloads. This model is a quantization of MaziyarPanahi/ NSFW_DPO_Noromaid-7b-Mistral-7B-Instruct-v0.1, which is also dual licensed in the same way. The model card forMaziyarPanahi/NSFW_DPO_Noromaid-7b-Mistral-7B-Instruct-v0.1 describes it as a merge of two additional models. The first, mistralai/Mistral-7B-Instruct-v0.1, is under the apache-2.0 license, and the second, athirdpath/NSFW_DPO_Noromaid-7b, is under the cc-by-nc-4.0 license. Following the supply chain back further, we observe that the dataset that athirdpath/ NSFW_DPO_Noromaid-7b was trained on was also licensed under cc-by-nc-4.0. mistralai/Mistral-7B-Instruct-v0.1 is a finetune of mistralai/Mistral-7B-v0.1, but there is no record of the data used for training. From all of this we see that the licensing relationship in this multi-licensing example is likely intended to be an AND (i.e., that the model is released under the terms of both the apache-2.0 and cc-by-nc-4.0 licenses). However, if other model trainers were to interpret the relationship as an OR (i.e., the model is licensed such that creators of derivative works can freely choose which licensing terms, either the apache-2.0 or cc-by-nc-4.0 license, to follow), it is possible that the non-commercial limitation imposed by the cc-by-nc-4.0 license could be dropped and non-compliance could be introduced. We also note that here, instead of finding a single suitable license for MaziyarPanahi/NSFW_DPO_Noromaid-7b-Mistral-7B-Instruct-v0.1, the model owner opted to just include the licenses of both dependencies. While not observed in this case, it is possible that this strategy could fail in certain instances where dependency licenses conflict.
@@
Some licenses, like Llama3, impose a generational limitation (Laurent, 2004) that specifies naming requirements for derivative works. According to the license, models that are built on a model under this license must have names prefixed with “llama3” (Meta, 2024). In our dataset, there are 3,335 models licensed under llama3. However, of these, only 473 (14.2%) models correctly have names beginning with “llama3” as required by the license. Only 768 (23%) models even contain the term “llama3” in their name. Indeed, just 2,421 (72.6%) models contain the term “llama” at all, and 384 (11.5%) only contain the string “l3.” The problem is likely even worse in practice, as these statistics only include models that documented that they were under the llama3 license.